OpenStudy (marziajackson$):

Which statement is true concerning the graph of two lines showing the given rates? Rate 1: 20 ft in 5 s Rate 2: 18 ft in 3 s A. The graph with Rate 1 passes through (3, 12). B. The graph with Rate 2 passes through (3, 15). C. The graph with Rate 1 passes through (8, 2). D. The graph with Rate 2 passes through (12, 2).

OpenStudy (marziajackson$):

@mathsciencehistory can you help me please?

OpenStudy (anonymous):

i might be able to

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

its c do you want to know how or do you just want the answer

OpenStudy (marziajackson$):

you can tell me how

OpenStudy (anonymous):

The equation for the line is y = slope * x For line 1: The rate or slope is 20 ft for 5s i.e 4ft/s For line 2: The rate is 6ft/s Now write the equations and see for the options. Hope this helps.

OpenStudy (anonymous):

Ok So line 1 is y = 4*x and line 2 is y = 6*x Put (3,12) and see if this satisfies any equation.

OpenStudy (anonymous):

im just using someone elses replies the first part is it

OpenStudy (anonymous):

the second is is more explanation if you dont get the first part